# Simulation of Polynomial-Time Nondeterministic Turing Machines via Universal Nondeterministic Turing Machines

### v18

@@ 7. Rich Structure of ${\rm co}\mathcal{NP}$
Our main goal in this section is to establish a result saying that similar to the complexity class $\mathcal{NP}$ having a rich structure, the complexity class ${\rm co}\mathcal{NP}$ also has intermediate languages that are not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete. To do so in a simple way, we need to quote the following useful result whose proof can be found in [Lad75]:
@@
It is sufficient to show that for every language $L\in{\rm co}\mathcal{NP}$, $L\leq_{p}{\rm TAUT}$ where $\leq_{p}$ is polynomial-time many-one reduction [Kar72]. We can modify the Cook-Levin reduction [Coo71, Lev73] from $\overline{L}=\{0,1\}^{*}-L$ to ${\rm SAT}$ (recall that all nondeterministic Turing machines are with input alphabet $\{0,1\}$ because that will be all we need, see Section 3). So for any input $w\in\{0,1\}^{*}$ the Cook-Levin reduction produces a formula $\varphi_{w}$ that is satisfiable if and only if $w\in\overline{L}$. In other words, the formula $\neg\varphi_{x}$ is in ${\rm TAUT}$ if and only if $x\in L$, which completes the proof.
@@
Then $L\in{\rm co}\mathcal{NP}$, and we will show that $L$ is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete. To see so, suppose that $L\in\mathcal{P}$; then since $\mathcal{P}={\rm co}\mathcal{P}$, we have $\overline{L}=L_{\rm inter}\in\mathcal{P}$, a contradiction to the fact that $L_{\rm inter}\notin\mathcal{P}$ by Lemma 7.1. Further assume that $L$ is ${\rm co}\mathcal{NP}$-complete, then ${\rm TAUT}\leq_{p}L$. By Lemma 7.2, $\overline{L}=L_{\rm inter}$ is $\mathcal{NP}$-complete; it is also a contradiction to the fact that $L_{\rm inter}$ is not $\mathcal{NP}$-complete. Thereby, to summarize from the above two cases, we have shown that $L$ is ${\rm co}\mathcal{NP}$-intermediate. This finishes the proof.
@@ 8. Frege Systems
The focal point of this section is Frege systems, which are very strong proof systems in a propositional setting [CR79], based on axiom schemes and rules such as modus ponens. While Frege systems operate with Boolean formulas as lines, the extended Frege system EF works with Boolean circuits; see e.g. [Jer05]. Furthermore, showing lower bounds on Frege or even extended Frege systems constitutes a major open problem in proof complexity; see e.g. [BP01].
Informally, a proof system for a language $\mathcal{L}$ is a definition of what is considered to be a proof that $\Phi\in\mathcal{L}$; see e.g. [CR79]. The key features of a proof system are that it is sound, i.e., only formulas in $\mathcal{L}$ have proofs, and complete, i.e., all formulas in $\mathcal{L}$ have proofs, and that there is an algorithm with running time polynomial in $|\pi|$ to check whether $\pi$ is a proof that $\Phi\in\mathcal{L}$ [BH19].
@@
If $L\subseteq\Sigma^{*}$, a proof system for $L$ is a function $f:\Sigma_{1}^{*}\rightarrow L$ for some alphabet $\Sigma_{1}$ and $f\in\mathfrak{F}$ such that $f$ is onto. We say that the proof system is polynomially bounded iff there is a polynomial $p(n)$ such that for all $y\in L$ there is a $x\in\Sigma_{1}^{*}$ such that $y=f(x)$ and $|x|\leq p(|y|)$, where $|z|$ denotes the length of a string $z$.
If $y=f(x)$, then $x$ is a proof of $y$, and $x$ is a short proof of $y$, if in addition, $|x|\leq p(|y|)$. Thus a proof system $f$ is polynomially bounded iff there is a bounding polynomial $p(n)$ with respect to which every $y\in L$ has a short proof.
In particular, Frege proof systems are proof systems for propositional logic. As a matter of fact, Frege systems are the usual “textbook” proof systems for propositional logic based on axioms and rules; see e.g. [CR79]. A Frege system composed of a finite set of axiom schemes and a finite number of rules is a possible axiom scheme. Furthermore, a Frege proof is a sequence of formulas where each formula is either a substitution instance of an axiom or can be inferred from previous formulas by a valid inference rule. At the same time, Frege systems are required to be sound and complete. However, the exact choice of the axiom schemes and rules does not matter, as any to Frege system are polynomial equivalent; see e.g.[CR79] or see Theorem 8.1 below. Thus, we can assume without loss of generality that modus ponens (see e.g. Remark 8.3 below) is the only rule of inference.
@@ A Frege proof system $F$ is an inference system for propositional logic based on
We give a more intuitive explanation about the inference rule of Modus Ponens: For example, in the proof of Theorem 5.2 in Section 5 where we prove that $L_{s}\in\mathcal{NP}$, $A$ is the following:
@@
If $f_{1}:\Sigma_{1}^{*}\rightarrow L$ and $f_{2}:\Sigma_{2}^{*}\rightarrow L$ are proof system for $L$, then $f_{2}$ $p$-simulates $f_{1}$ provided there is a function $g:\Sigma_{1}^{*}\rightarrow\Sigma_{2}^{*}$ such that $g\in\mathfrak{F}$, and $f_{2}(g(x))=f_{1}(x)$ for all $x$.
@@ 8.1. Lower Bounds for Frege Systems
Let $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ be two arbitrary Frege systems, the following theorem indicates that these standard proof systems for the propositional calculus are about equally powerful:
@@
Any two Frege systems $p$-simulate each other. Hence one Frege system is polynomially bounded if and only if all Frege systems are.
@@
The “only if” part. Suppose that $\mathcal{NP}$ is closed under complementation. The the complement of the set of tautologies is in $\mathcal{NP}$, since to verify that a formula is not a tautology one can guess at a truth assignment and verify that it falsifies the formula.
Since nondeterministic Turing machines can simulate Frege proof (see e.g. [Bus99]) and by Theorem 8.1 we know that any two Frege systems $p$-simulate each others, we then can present the proof of Theorem 1.8 as follows:
Proof of Theorem 1.8. We show the theorem by contradiction. Suppose that there is a polynomial $p(n)$ such that for all $\psi\in TAUT$, there exists a proof $\oint$ of $\psi$ of length at most $p(|\psi|)$ (see Definition 8.4), then a nondeterministic Turing machine $M$ on input $\psi$ can guess $\oint$ of length at most $p(|\psi|)$ and check whether $\oint$ is correct in deterministic polynomial time, and we know that the set of tautologies is ${\rm co}\mathcal{NP}$-complete, which, together with Theorem 8.2, implies
@@ 9. Open Problems
Does $\mathcal{P}=\mathcal{NP}\cap{\rm co}\mathcal{NP}$? Currently, we have no answer to this problem. We tried to think about it but failed. We remark that this is a fascinating open problem in complexity theory. It is easy to see that $\mathcal{P}\subseteq\mathcal{NP}\cap{\rm co}\mathcal{NP}$ and the real problem is whether $\mathcal{NP}\cap{\rm co}\mathcal{NP}\subseteq\mathcal{P}$. Note that an example of a problem that is known to belong to both $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ (but not known to be in $\mathcal{P}$) is Integer factorization [RSA78]: given positive integers $m$ and $n$, determine if $m$ has a factor less than $n$ and greater than one.
In addition, does the complexity class $\mathcal{BQP}$ (see page 213 of [AB09]) contains the problem Satisfiability? Note that by footnote $1$, we have $\mathcal{BQP}\neq\mathcal{NP}$, but this can not exclude that $\mathcal{NP}\subsetneqq\mathcal{BQP}$. If one can show that Satisfiability is not in $\mathcal{BQP}$, then the relationship between $\mathcal{BQP}$ and $\mathcal{NP}$ were clear. Currently, we have no answer to this problem. We tried to think about it but also failed. We remark that this is also a fascinating open problem in complexity theory.
@@ Appendix A Two Definitions of $\mathcal{NP}$ Are Equivalent
The official descriptions on the definitions of $\mathcal{NP}$ in [Coo00] is given by:
@@
$\mathcal{NP}\overset{{\rm def}}{=}\{L\,:\,w\in L\Leftrightarrow\exists y(|y|\leq|w|^{k}\,\,\,\text{and }R(w,y))\}$
@@
$T_{M}(n)\leq n^{k}+k,\quad\forall n\in\mathbb{N}_{1}$
@@
As you may note that, this definition (i.e., Definition A.1) looks a little different from that given in Section 2. But they are equivalent. Here, we will show the equivalence for these two definitions of $\mathcal{NP}$.
@@
To begin, we should first give another definition of $\mathcal{NP}$ which is similar to the official descriptions of the definition of $\mathcal{P}$ in [Coo00] (i.e., Definition A.2 above):
@@
We need to show that Definition A.3 is equivalent to Definition A.1 and to show the equivalence between Definition A.3 and the definition of $\mathcal{NP}$ given in Section 2. By a simple argument, it is clear that Definition A.3 and the definition of $\mathcal{NP}$ given in Section 2 are equivalent. We proceed as follows:
@@
$T_{M}(|w|)\leq c_{0}n^{k}+c_{1}n^{k-1}+\cdots+c_{k-1}n+c_{k}\quad\text{where $c_{0}>0$}.$
@@
For any constants $c_{0}$, $c_{1}$, $\cdots$, $c_{k}$, there must exists a minimal $t\in\mathbb{N}_{1}$ such that for all $n\in\mathbb{N}$ and for any $w\in\Sigma^{n}$
@@
It is time for us to show that Definition A.3 is equivalent to Definition A.1. For the purposes of brevity, we omit the proof and refer the reader to [Kar72] (cf. Theorem 1 in [Kar72]), or to the proof of Theorem $7$.20202020 in [Sip13] which is a theorem answering the same question discussed here (see p. $294$ of [Sip13]). Thus this completes the proof.

### v19

@@ 7. Rich Structure of ${\rm co}\mathcal{NP}$
Our main goal in this section is to establish a result saying that, similar to the complexity class $\mathcal{NP}$ having a rich structure, the complexity class ${\rm co}\mathcal{NP}$ also has intermediate languages that are not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete. To do so in a simple way, we need to quote the following useful result whose proof can be found in [Lad75]:
@@
It is sufficient to show that for every language $L\in{\rm co}\mathcal{NP}$, $L\leq_{p}{\rm TAUT}$, where $\leq_{p}$ is polynomial-time many-one reduction [Kar72]. We can modify the Cook-Levin reduction [Coo71, Lev73] from $\overline{L}=\{0,1\}^{*}-L$ to ${\rm SAT}$ (recall that all nondeterministic Turing machines are with input alphabet $\{0,1\}$ because that will be all we need; see Section 3). So for any input $w\in\{0,1\}^{*}$, the Cook-Levin reduction produces a formula $\varphi_{w}$ that is satisfiable if and only if $w\in\overline{L}$. In other words, the formula $\neg\varphi_{x}$ is in ${\rm TAUT}$ if and only if $x\in L$, which completes the proof.
@@
Then $L\in{\rm co}\mathcal{NP}$, and we will show that $L$ is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete. To see so, suppose that $L\in\mathcal{P}$; then since $\mathcal{P}={\rm co}\mathcal{P}$, we have $\overline{L}=L_{\rm inter}\in\mathcal{P}$, a contradiction to the fact that $L_{\rm inter}\notin\mathcal{P}$ by Lemma 7.1. Further assume that $L$ is ${\rm co}\mathcal{NP}$-complete; then ${\rm TAUT}\leq_{p}L$. By Lemma 7.2, $\overline{L}=L_{\rm inter}$ is $\mathcal{NP}$-complete; it is also a contradiction to the fact that $L_{\rm inter}$ is not $\mathcal{NP}$-complete. Thereby, to summarize from the above two cases, we have shown that $L$ is ${\rm co}\mathcal{NP}$-intermediate. This finishes the proof.
@@ 8. Frege Systems
The focal point of this section is Frege systems, which are very strong proof systems in a propositional setting [CR79], based on axiom schemes and rules such as modus ponens. While Frege systems operate with Boolean formulas as lines, the extended Frege system EF works with Boolean circuits; see e.g., [Jer05]. Furthermore, showing lower bounds on Frege or even extended Frege systems constitutes a major open problem in proof complexity; see e.g., [BP01].
Informally, a proof system for a language $\mathcal{L}$ is a definition of what is considered to be a proof that $\Phi\in\mathcal{L}$; see e.g., [CR79]. The key features of a proof system are that it is sound, i.e., only formulas in $\mathcal{L}$ have proofs, and complete, i.e., all formulas in $\mathcal{L}$ have proofs, and that there is an algorithm with running time polynomial in $|\pi|$ to check whether $\pi$ is a proof that $\Phi\in\mathcal{L}$ [BH19].
@@
If $L\subseteq\Sigma^{*}$, a proof system for $L$ is a function $f:\Sigma_{1}^{*}\rightarrow L$ for some alphabet $\Sigma_{1}$ and $f\in\mathfrak{F}$ such that $f$ is onto. We say that the proof system is polynomially bounded iff there is a polynomial $p(n)$ such that for all $y\in L$ there is an $x\in\Sigma_{1}^{*}$ such that $y=f(x)$ and $|x|\leq p(|y|)$, where $|z|$ denotes the length of a string $z$.
If $y=f(x)$, then $x$ is a proof of $y$, and $x$ is a short proof of $y$ if, in addition, $|x|\leq p(|y|)$. Thus a proof system $f$ is polynomially bounded iff there is a bounding polynomial $p(n)$ with respect to which every $y\in L$ has a short proof.
In particular, Frege proof systems are proof systems for propositional logic. As a matter of fact, Frege systems are the usual‘ ‘textbook” proof systems for propositional logic based on axioms and rules; see e.g., [CR79]. A Frege system composed of a finite set of axiom schemes and a finite number of rules is a possible axiom scheme. Furthermore, a Frege proof is a sequence of formulas where each formula is either a substitution instance of an axiom or can be inferred from previous formulas by a valid inference rule. At the same time, Frege systems are required to be sound and complete. However, the exact choice of the axiom schemes and rules does not matter, as any Frege systems are polynomially equivalent; see e.g., [CR79] or see Theorem 8.1 below. Thus, we can assume without loss of generality that modus ponens (see e.g., Remark 8.3 below) is the only rule of inference.
@@ A Frege proof system $F$ is an inference system for propositional logic based on
We give a more intuitive explanation about the inference rule of Modus Ponens: For example, in the proof of Theorem 5.2 in Section 5, where we prove that $L_{s}\in\mathcal{NP}$, $A$ is the following:
@@
If $f_{1}:\Sigma_{1}^{*}\rightarrow L$ and $f_{2}:\Sigma_{2}^{*}\rightarrow L$ are proof systems for $L$, then $f_{2}$ $p$-simulates $f_{1}$ provided there is a function $g:\Sigma_{1}^{*}\rightarrow\Sigma_{2}^{*}$ such that $g\in\mathfrak{F}$, and $f_{2}(g(x))=f_{1}(x)$ for all $x$.
@@ 8.1. Lower Bounds for Frege Systems
Let $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ be two arbitrary Frege systems; the following theorem indicates that these standard proof systems for the propositional calculus are about equally powerful:
@@
Any two Frege systems $p$-simulate each other. Hence, one Frege system is polynomially bounded if and only if all Frege systems are.
@@
The “only if” part. Suppose that $\mathcal{NP}$ is closed under complementation. Then the complement of the set of tautologies is in $\mathcal{NP}$, since to verify that a formula is not a tautology, one can guess at a truth assignment and verify that it falsifies the formula.
Since nondeterministic Turing machines can simulate Frege proofs (see e.g., [Bus99]), and by Theorem 8.1 we know that any two Frege systems $p$-simulate each other, we can then present the proof of Theorem 1.8 as follows:
Proof of Theorem 1.8. We show the theorem by contradiction. Suppose that there is a polynomial $p(n)$ such that for all $\psi\in TAUT$, there exists a proof $\oint$ of $\psi$ of length at most $p(|\psi|)$ (see Definition 8.4); then a nondeterministic Turing machine $M$ on input $\psi$ can guess $\oint$ of length at most $p(|\psi|)$ and check whether $\oint$ is correct in deterministic polynomial time, and we know that the set of tautologies is ${\rm co}\mathcal{NP}$-complete, which, together with Theorem 8.2, implies
@@ 9. Open Problems
Does $\mathcal{P}=\mathcal{NP}\cap{\rm co}\mathcal{NP}$? Currently, we have no answer to this problem. We tried to think about it but failed. We remark that this is a fascinating open problem in complexity theory. It is easy to see that $\mathcal{P}\subseteq\mathcal{NP}\cap{\rm co}\mathcal{NP}$, and the real problem is whether $\mathcal{NP}\cap{\rm co}\mathcal{NP}\subseteq\mathcal{P}$. Note that an example of a problem that is known to belong to both $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ (but not known to be in $\mathcal{P}$) is Integer factorization [RSA78]: given positive integers $m$ and $n$, determine if $m$ has a factor less than $n$ and greater than one.
In addition, does the complexity class $\mathcal{BQP}$ (see page 213 of [AB09]) contain the problem Satisfiability? Note that by footnote $1$, we have $\mathcal{BQP}\neq\mathcal{NP}$, but this can not exclude that $\mathcal{NP}\subsetneqq\mathcal{BQP}$. If one can show that the problem of Satisfiability is not in $\mathcal{BQP}$, then the relationship between $\mathcal{BQP}$ and $\mathcal{NP}$ will were clear. Currently, we have no additional information about this problem. We tried to think about it but also failed. We remark that this is also a fascinating open problem in complexity theory.
@@ Appendix A Two Definitions of $\mathcal{NP}$ Are Equivalent
The official description of the definition of $\mathcal{NP}$ in [Coo00] is given by the following.
@@
$\mathcal{NP}\overset{{\rm def}}{=}\{L\,:\,w\in L\Leftrightarrow\exists y(|y|\leq|w|^{k}\,\,\,\text{and }R(w,y))\},$
@@
$T_{M}(n)\leq n^{k}+k,\quad\forall n\in\mathbb{N}_{1},$
@@
As you may note, this definition (i.e., Definition A.1) looks a little different from that given in Section 2. But they are equivalent. Here, we will show the equivalence for these two definitions of $\mathcal{NP}$.
@@
To begin, we should first give another definition of $\mathcal{NP}$, which is similar to the official descriptions of the definition of $\mathcal{P}$ in [Coo00] (i.e., Definition A.2 above):
@@
We need to show that Definition A.3 is equivalent to Definition A.1 and to show the equivalence between Definition A.3 and the definition of $\mathcal{NP}$ given in Section 2.
By a simple argument, it is clear that Definition A.3 and the definition of $\mathcal{NP}$ given in Section 2 are equivalent. We proceed as follows:
@@
$T_{M}(|w|)\leq c_{0}n^{k}+c_{1}n^{k-1}+\cdots+c_{k-1}n+c_{k}\quad\text{where $c_{0}>0$},$
@@
For such constants $c_{0}$, $c_{1}$, $\cdots$, $c_{k}$, there must exist a minimal $t\in\mathbb{N}_{1}$ such that for all $n\in\mathbb{N}$ and for any $w\in\Sigma^{n},$ we have that
@@
It is time for us to show that Definition A.3 is equivalent to Definition A.1. For the purposes of brevity, we omit the proof and refer the reader to [Kar72] (cf. Theorem 1 in [Kar72]), or to the proof of Theorem $7$.20202020 in [Sip13], which is a theorem answering the same question discussed here (see p. $294$ of [Sip13]). Thus, this completes the proof.